OpenStudy (anonymous):

A 50 loop coil of radius 3.0 cm is oriented so that its normal is parallel to an external field B which increases from 0.1 T to .35 T in 2.0 milliseconds. Find the average emf in the coil.

OpenStudy (anonymous):

\[EMF=(d \phi/dt)\] \[=A \times(dB/dt)\] \[=\Pi \times r ^{2} \times((0.35-0.1)\div(2\times10^{-3}))\] EMF= 0.35 volt